OpenStudy (codysek98):

I need explanations on how to take the derivative of a trig functions. Could somebody explain how to do so and then help me work out a few examples?

OpenStudy (dumbcow):

Sure, with basic trig functions. Learning the derivatives is simple memorization or using a cheat sheet. \[\frac{d}{dx} \sin x = \cos x\] \[\frac{d}{dx} \cos x = -\sin x\] \[\frac{d}{dx} \tan x = \sec^2 x\]

OpenStudy (codysek98):

And for csc, sec, cot?

OpenStudy (dumbcow):

they can be found by writing in terms of sin,cos,or tan , Then use the chain rule. Anyway here are the derivatives....you can work them out on your ow. \[\frac{d}{dx} \csc x = -\csc x \cot x\] \[\frac{d}{dx} \sec x = \sec x \tan x\] \[\frac{d}{dx} \cot x = -\csc^2 x\]

OpenStudy (codysek98):

So, if I'm taking the derivative of a trig function, do I basically just plug that in in place of it? So if I need to take the derivative of lets say, 3cosx, it would just be -3sinx?

OpenStudy (dumbcow):

correct, thats it.. However if its cos(3x), you must use chain rule ----> -3sin(3x)

OpenStudy (codysek98):

Can you explain the chain rule to me a bit more in detail? Someone helped me by using it in one of my last questions, but I don't *fully* understand it. I know you multiply whatever is inside of the parenthesis at the end (i don't know how to word it). So in any situation, if something is being multiplied by something in parenthesis, do I just multiply it at the end in the derivative?

OpenStudy (dumbcow):

yes thats pretty much the right idea. lets do an example \[\sin^2 x\] So you want to think of 2 functions, theres the sin function and the squared function. \[ u = \sin x \rightarrow f(u) = u^2\] chain rule says: \[f'(u) = 2u * du\] where du is derivative of "u" function, in this case sin x du = cos x \[f'(u) = 2u*du = 2\sin x \cos x\]

OpenStudy (codysek98):

Okay, and will the chain rule ALWAYS say 2u*du? Or is that specific for this problem? Would it be 3u*du if the example was sin^3(x)?

OpenStudy (dumbcow):

Another one: \[\sin(3x^2 -1)\] In this case, let \[ u = 3x^2 - 1 \rightarrow f(u) = \sin u\] \[f'(u) = \cos u * du\] \[du = 6x\] \[f'(u) = \cos(3x^2 -1) (6x)\]

OpenStudy (dumbcow):

It depends on the function. you have the right idea for sin^3, however it would be 3u^2 du

OpenStudy (codysek98):

It's making a lot more sense. Now, how do I determine which part of the function is u and which part is f(u)? Is u always the part of the function that is being altered? Like, in the first example, u was the term that was being squared.

OpenStudy (dumbcow):

right thats the trick . I like to think of the "u" function as being the inner function While f(u) is the outer base function

OpenStudy (dumbcow):

\[\frac{1}{x^3 -1}\] what is u?

OpenStudy (codysek98):

u= x^3-1 and f(u) = 1/u Is that right?

OpenStudy (dumbcow):

yes, good job

OpenStudy (dumbcow):

sorry did you have your own questions?

OpenStudy (codysek98):

Yes, I have a couple if you have the time to help me through them.

OpenStudy (dumbcow):

sure

OpenStudy (codysek98):

Okay. So this is one of the simpler ones. y=8xsinx So, I know the derivative of sinx is cosx.

OpenStudy (codysek98):

Would u be sinx and f(u) be 8x(u)?

OpenStudy (dumbcow):

note, f(u) cannot have an "x" in it This looks more like a problem for product rule

OpenStudy (dumbcow):

the u, f(u) notation only works where chain rule applies

OpenStudy (codysek98):

Is it just a matter of recognizing if chain rule would work for a function or is there some sort of process I can do to check if the function would require chain rule?

OpenStudy (dumbcow):

yep, first check to see if you only have base functions or if its a more complicated function. in this case you just have 2 base functions multiplied together

OpenStudy (codysek98):

What's the product rule? I know the answer to the equation is 16cosx-8xsinx

OpenStudy (dumbcow):

product rule \[(fg)' = f'g + fg'\] f = 8x, g = sin x

OpenStudy (dumbcow):

your answer is wrong, are you sure its 8x*sin(x) ?

OpenStudy (codysek98):

OH wait! y' means the first derivative, and y'' means second derivative, correct?

OpenStudy (codysek98):

If so, then I'm needing to find the second derivative.

OpenStudy (codysek98):

Of y=8xsinx

OpenStudy (dumbcow):

ahh, yes y'' means 2nd derivative

OpenStudy (codysek98):

Sorry about that! So to find second derivatives, is it easier to find the first derivative and then take the derivative of that? Or is there another process that I should be doing?

OpenStudy (dumbcow):

unfortunately there's no shortcut. You have to find y' first . Then take derivative again

OpenStudy (codysek98):

Okay. So for the first derivative, it would be 8sinx+8x(cosx), right?

OpenStudy (dumbcow):

correct

OpenStudy (codysek98):

So how do I go about finding the derivative of that? Do I have to simplify 8x(cosx) if possible?

OpenStudy (codysek98):

The derivative of 8sinx would be 8cosx, right? And then the derivative of 8x(cosx)...would just be 8xsinx?

OpenStudy (dumbcow):

you take derivative of each term You will have to use product rule again for the "8xcosx" term

OpenStudy (codysek98):

Okay, so the derivative of 8sinx is 8cosx. And the derivative of "8xcosx" would be 8(sinx)+8x(sinx) ?

OpenStudy (codysek98):

8(cosx)+8x(sinx) **

OpenStudy (dumbcow):

almost, remember derivative of cos = -sin

OpenStudy (codysek98):

ohhh, that's right. So 8(cosx)+8x(-sinx) ?

OpenStudy (codysek98):

So in all, I would have 8cosx+8cosx+8x(-sinx) ? The two "8cosx" would add to 16cosx+8x(-sinx). And then how do I go further than that?

OpenStudy (dumbcow):

you got it. you cant simplify it any further

OpenStudy (codysek98):

It's saying the correct answer is 16cosx-8xsinx though, would I need to factor out the - from the sinx?

OpenStudy (dumbcow):

yes

OpenStudy (codysek98):

So it's done! Awesome! Do you want to help me out on one of the more complicated ones? (At least they seem more complicated to me)

OpenStudy (dumbcow):

sure

OpenStudy (codysek98):

Find dy/dt. y=cos^6(pi(t)-9)

OpenStudy (codysek98):

\[y=\cos^6(\pi(t)-9)\])

OpenStudy (codysek98):

First, is there a difference between finding y' and finding dy/dt?

OpenStudy (dumbcow):

no they are the same thing. So this is a nested function. We need to use chain rule twice.

OpenStudy (codysek98):

Okay, so I'm guessing I do the chain rule in the parenthesis first. The pi(t)-9

OpenStudy (codysek98):

And u would be...maybe pi(t) and f(u) = u-9?

OpenStudy (dumbcow):

not quite, that makes f(u) too simple try u = pi(t) - 9

OpenStudy (codysek98):

So f(u) = cos^6(u)?

OpenStudy (dumbcow):

yes

OpenStudy (codysek98):

Okay. And the derivative of pi(t)-9 would be pi and the derivative of cosu would be -sinu...so it would be -sin(pi) ?

OpenStudy (codysek98):

wait no it would be

OpenStudy (codysek98):

-6sin^5(pi) ?

OpenStudy (codysek98):

-6sin^5(pi(t)-9)(pi) ?

OpenStudy (dumbcow):

getting closer :) Important when dealing with trig functions, the derivative will ALWAYS keep the same input in parenthesis as original In this case --->( pi(t) - 9) The derivative du = pi is just multiplied on the outside. Ok this is where i said we need to do chain rule twice \[\cos^6 u\] \[w = \cos u \rightarrow f(w) = w^6\] \[f'(w) = 6w^5 * dw*du\] \[= 6 \cos^5 u (-\sin u) (\pi)\]

OpenStudy (dumbcow):

Final step is plug back in (pi)t - 9 for u

OpenStudy (codysek98):

And I can factor out that -1, so I could have -6(pi)cos^5(pi(t)-9)(sin(pi(t)-9)

OpenStudy (codysek98):

Let me write it in the equation thing, one sec.

OpenStudy (codysek98):

\[-6(\pi)\cos^5((\pi)(t)-9)(\sin((\pi)(t)-9)\]

OpenStudy (dumbcow):

very good This is called a nested function where you have to use chain rule multiple times. It can take getting used too

OpenStudy (codysek98):

And how did you determine that it was a nested function? Is it because the cos had an exponent? So if it was just cos(pi(t)-9), it wouldn't be a nested function?

OpenStudy (dumbcow):

correct

OpenStudy (codysek98):

Jesus, this is a lot to take in!

OpenStudy (dumbcow):

yes, welcome to the world of calculus. But hopefully you have weeks to learn, is this h.s or college?

OpenStudy (codysek98):

College.

OpenStudy (codysek98):

Would you like to explain implicit differentiation? I have a few examples of that I need to complete.

OpenStudy (dumbcow):

sure, is this an online class? you seem to be covering multiple sections at once

OpenStudy (codysek98):

Yes it is. I'm covering many, many sections at once.

OpenStudy (dumbcow):

thats typically not the best way to learn math , oh well

OpenStudy (codysek98):

For the upcoming test, I have everything that you've covered with me so far. As well as Continuous/differentiable points on a graph, finding speed/acceleration/velocity of an equation, graphing derivatives, implicit differentiation, tangents to a curve, slopes of a curve, how fast a wheel is rolling...how fast water is being drained from a container..

OpenStudy (codysek98):

My professor kind of dumps like 12-15 assignments online and then gives us a due date for them and then announces the test later along the line.

OpenStudy (dumbcow):

ok well for implicit, you need to know the product rule really well Also just remember that only the "y" term derivative will be multiplied by dy/dx Example: \[xy - x^2 = 1\] Take derivative of each term \[(xy)' = (1)y + x(1)\frac{dy}{dx}\] \[(-x^2)' = -2x\] \[1' = 0\] \[\rightarrow (y +x \frac{dy}{dx}) -2x = 0\] Then you have to algebraically solve for "dy/dx" \[\frac{dy}{dx} = \frac{2x-y}{x}\]

OpenStudy (dumbcow):

heres a good website with lots of notes that will help you. http://tutorial.math.lamar.edu/Classes/CalcI/CalcI.aspx